OpenStudy (anonymous):

Find an equation of the line perpendicular to the graph of 14x – 7y = 1 that passes through the point at (–2, 4)

OpenStudy (ivettef365):

first convert the equation in form y=mx+b, to find the slope (m)

OpenStudy (anonymous):

Yes - find that slope of the equation given. Then, since your line is supposed to be perpendicular, you won't use that slope, but you WILL use the opposite reciprocal of that slope. So, for instance, if you found the slope of that line was 2/3, you would use -3/2 for your slope (opposite sign AND reciprocal). So you use your slope that you found and the points given in the problem in the Point-Slope formula. POINT-SLOPE FORMULA: \[y-y _{1}=m(x-x _{1})\]